# Modeling wildfire dynamics through a physics-based approach incorporating fuel moisture and landscape heterogeneity

### v1

@@ 4.5 Two-dimensional domains
We now extend the spatial domain to two spatial dimensions (2D) and consider different scenarios with spatial heterogeneity. Below, results are presented for four different cases and the observations made above for one spatial dimension are now extended to more realistic fire propagation patterns in two-dimensional landscapes.
@@ 4.5.1 Propagation in a homogeneous environment: assessing the effect of fuel moisture
The aim of this case is to assess the effect of fuel moisture in 2D fire propagation across homogeneous land, under a low wind velocity $\mathbf{w}=(0.25,0.25)^{T}$ m/s. We consider a domain of $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
with $(x_{1},y_{1})=(100,100)$ m. Under these wind conditions, we expect the predominant propagation direction to point toward the top-right corner of the domain.
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 1400 seconds using $\mathsf{CFL}=0.4$. We consider the complete model with radiation and constant reaction rate, assuming the simplification in Eq. (9). The S2M model is used to represent the effect of moisture. The model parameters are taken from Tab. 24, unless otherwise specified.
We compare four different simulations considering different moisture levels, $M=0.0$, $M=0.2$, $M=0.4$, $M=0.6$. Fig. 15 shows the predicted temperature and biomass distribution at the final time. Besides, the fire isochrones are depicted in red every 200 seconds. These results evidence that the fuel moisture level has a strong impact on the shape of the burned area, as well as on the rate of spread. When considering zero moisture, the propagation takes place in all directions, being slightly faster in the direction of wind. Contrarily, when increasing the moisture level, propagation mainly takes place in the direction of wind. Besides, the width of the fire front and the rate of spread reduces as the moisture level increases. From a geometric perspective, the angle of the fire front decreases with increasing moisture level.
@@
Figure 15: Fire spread for different moisture content levels with a wind velocity $\mathbf{w}=(0.25,0.25)^{T}$ m/s.
The reduction of the rate of spread in direction of the wind is in accordance with Fig. 11 showing the decrease in spreading rate with increasing moisture level. More detailed, we can see that the spread orthogonal to the wind direction is reduced the most. This is in line with the zero wind case of Fig. 11 where the rate of spread drops to zero for moisture levels above 80%. In contrast, for non-zero wind speed, the decrease in the rate of spread is less significant and reaches a non-zero threshold. The two-dimensional simulation combines these two results when regarding either the direction of the wind or the orthogonal direction. As the moisture levels regarded are all below the critical threshold of $M=0.8$, there is still a fire spread orthogonal to the wind velocity, but the rate of spread in this direction is small.
Tab. 7 shows the total burned area at the final time of the simulation. If the combustion process starts, the biomass reduces below the threshold $Y=0.99$ and we indicate the area to be burned. The results show a strong reduction of burned area for higher moisture levels, e.g. a reduction of 80% for a moisture level $M=0.2$ compared to the no moisture case.
@@ 4.5.2 Propagation in heterogeneous environments: assessing the effect of fuel moisture
The aim of this case is to understand the model behavior when changing the fuel moisture in a heterogeneous landscape composed of four quadrants with different fuel moisture content. We consider a domain of $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 1400 seconds using $\mathsf{CFL}=0.4$. We consider the complete model with radiation and constant reaction rate, assuming the simplification in Eq. (9). The S2M model is used to represent the effect of moisture. The model parameters are taken from Tabs. 24, unless otherwise specified. We consider zero wind conditions.
@@
Figure 16: Effect of fuel moisture in a heterogeneous setting without wind: The four quadrants have different but spatially constant fuel moisture content, and the fire spreads from $(100,100)$ m.
@@ 4.5.3 Propagation in heterogeneous environments: assessing the effect of surface coverage
The aim of this case is to understand the model behavior when changing the fuel load by modifying the surface coverage, $SC$, assuming that the fuel load is a limiting factor for the parameters chosen here –i.e., corresponding to the region of Fig 13 where the ROS increases with fuel load–. We expect that those areas with a lower value of $SC$ act as a firebreak. We consider a domain of $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 2200 seconds using $\mathsf{CFL}=0.4$. We consider the complete model with radiation and constant reaction rate, assuming the simplification in Eq. (9). The S2M model is used to represent the effect of moisture. The model parameters are taken from Tabs. 24, unless otherwise specified above.
We compare four different simulations considering different fuel surface coverages, $SC_{r}$, namely $SC_{r}=0.2$, $SC_{r}=0.4$, $SC_{r}=0.6$, $SC_{r}=0.8$, setting a wind velocity $\mathbf{w}=(1,1)^{T}$ m/s. Figure 17 shows the predicted temperature and biomass distribution at the final time. Besides, the fire isochrones are depicted in red every 200 seconds. These results confirm that higher fuel loads, regarded as higher $SC$, enhance fire propagation. Besides, there is a threshold at which the fire extinguishes and propagation stops, acting as a firebreak. In this example, the threshold appears for a value of $SC$ between 0.2 and 0.3. Fig. 17 shows a strong dependency on stopping the fire on the surface coverage but, in oppose to the influence of the moisture content, the rate of spread is less strongly affected by the reduced surface coverage. This tendency is supported by the one-dimensional simulation in Fig. 14, where the drop from a higher rate of spread for high surface coverage values to zero for small surface coverage values is very abrupt.
@@
Figure 17: Fire spread at the interface of a changing surface coverage, starting from $(50,50)$ m.
In the previous results, the burned area was represented by showing the distribution of $Y$ in space at the final simulation time. Recall that $Y$ represents the fuel mass fraction with respect to the initial mass, and therefore it is normalized to a range from 0 to 1. To adequately represent the actual evolution of the fuel, we show the spatial distribution of $W$ in Fig. 18(right) and compare it to $Y$ in Fig. 18(left), for a surface coverage $SC_{r}=0.4$. In this case, we set $W_{f,\mathrm{max}}=4$ kg/m2, provided $\bar{\rho}_{f}=400$ kg/m3. $W(x,y,0)$ is $1.6$ kg/m2 on the left side ($SC=1$) and 4 kg/m2 on the right side ($SC=0.4$).
Figure 18: Comparison of the reduction of fuel fraction $Y$(left) and fuel load $W$ (left). The surface coverage is reduced to $SC=0.4$ for $x>100$ m.
At the final simulation time, the average fuel load on the burned area on the left-hand side is around 0.18 kg/m2, whereas it is 0.35 kg/m2 on the burned area on the right-hand side. This result indicates that a less dense biomass may lead to a higher remaining biomass after the burning. One reason for this behavior is a difference in the maximal temperature for the different initial fuel loads. In the region with higher initial fuel load, the maximal temperature is higher, indicating a stronger combustion and leading to fewer remaining fuel.
@@ 4.5.4 Propagation in heterogeneous environments with topography
This case examines fire propagation across uneven vegetation, both with and without the influence of topography. We consider a domain of $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 2600 seconds using $\mathsf{CFL}=0.4$. We consider the complete model with radiation and constant reaction rate, assuming the simplification in Eq. (9). The bulk density and heat capacity are computed using the more realistic mixture approximation in Eq. (8). We assume that there is zero moisture and the wind is given by $\mathbf{w}=(0.0,0.5)^{T}$ m/s. The model parameters are taken from Tables 24, unless otherwise specified above.
Fig. 19 shows a tendency to spread to areas with higher fuel load. Additionally, we find fire breaks for small fuel loads, which is comparable to the situation in Fig. 17(a). The wind direction is parallel to the $y$-axis, but we see a strong propagation as well in the areas with high fuel load orthogonal to the wind direction.
Fig. 19 and Fig. 20 show the spread of fire in the same biomass configuration but now with additional uneven topography. The topography acts as an additional advective influence, compare Sec. 2.3. The topography in Eq. (82) has a gradient pointing towards $(x,y)=(200,200)$ m, therefore we find an additional advective influence towards the upper right-hand corner. Fig. 19 shows this influence by having a stronger tendency to spread towards this corner.
@@
Figure 19: Simulation of spatially heterogeneous fuel with wind $\mathbf{w}=(0.0,0.5)^{T}$ m/s, $M=0.0$, and the mixture approximation in Eq. (8). Left: Temperature $T$, right: Fuel load $W$.
The three-dimensional Fig. 20 shows how the fire spread follows the inclination of the topography, influenced additionally by the heterogeneous fuel load. The results in Fig. 20 show that the developed model is able to capture more complex and realistic scenarios including heterogeneity in the biomass and the advection term.
Figure 20: Fire spread on an inclined topography with spatially heterogeneous fuel load. The wind velocity is wind $\mathbf{w}=(0.0,0.5)^{T}$ m/s, and the moisture content is $M=0.0$.
@@ 5 Conclusions
We presented an ADR wildfire propagation model that accounts for the effect of fuel moisture, and also considers wind, local radiation, natural convection, and topography. Special attention has been given to the derivation of the governing equations from the fundamental conservation laws within an idealized vegetation stratum which is modelled as a two-phase porous flow. We have outlined all the assumptions used to derive the ADR wildfire propagation model in its most commonly used form, providing a better insight into the model’s limitations.
@@
The model includes different mechanisms for live and dead fuel by using the apparent calorific capacity method, which models the effect of thermal phase changes through an apparent (or effective) heat capacity. Three different moisture models—the S2M, C2M, and S3M—have been presented to represent the evaporation process and the constituents of fuel with different levels of detail. From a practical point of view, the results show that the 3-stage moisture model (S3M) does not improve the predictions of the simple 2-stage model (S2M). On the other hand, the use of a complete moisture model (C2M), which distinguishes between live and dead biomass, proves very useful, because it allows for the representation of the different physicochemical mechanisms involved in the process. While live fuel moisture content can be estimated using remote sensing techniques and set as input data, dead fuel moisture content can be approximated from the local environmental conditions, e.g., relative humidity and ambient temperature. The approximation of dead fuel moisture in this way could significantly improve wildfire propagation predictions, because dead fuel moisture is often considered to be the main control on fire characteristics and fire impact on fuel and vegetation, see, e.g., [41, 60]. A limitation of this model is that we do not consider the variation in time of the relative humidity. A more complete approach, see e.g., [38], could be easily adapted. Also note that the utilization of dead fuel moisture data in the model, if available, is straightforward by substituting this approach –i.e., $M_{d}$ would be considered as an input parameter–.
Our work provides new insights on ADR-based wildfire propagation modeling that complement previous studies, e.g., [7, 26, 28, 34, 37, 42, 65], as follows. First, we show that the commonly used Arrhenius combustion function can be approximated by a constant as long as the switching function for activating and deactivating the combustion process is still included. This approximation is in line with the macroscopic scale of the wildfire model: The Arrhenius law is based on chemical processes on a much smaller length scale, and the resulting nonlinear combustion term contrasts the model accuracy on the macroscopic scale. Further, we show the relevance of the radiation leading to a nonlinear diffusion term. In combination with the moisture model, the nonlinear radiation leads to a more realistic value for the surviving biomass than previous models, where parameters leading to high temperatures, e.g. high wind velocities, were connected with large amounts of surviving biomass. The model including now moisture and radiation (see Fig. 8) shows a higher rate of spread with higher maximal temperatures for lower moisture levels, but the surviving biomass is similar.
Combustion and radiation are crucial for modeling wildfire propagation through advection–diffusion–reaction-type models. Based on our results, we recommend having a lower level of detail for the combustion function but a higher level of detail for the diffusion term including the nonlinear effects by radiation. The diffusion mechanism’s dependence on $T^{3}$ links the ROS to the medium’s energy, which is proportional to its temperature, ensuring physically consistent results and reproducing observed trends, see Rem. 7.
@@
The parametric analyses presented herein provide evidence of the model’s ability to reproduce the expected fire propagation behavior and the dependence of the ROS on the main drivers of fire spread. The model reproduces both the exponentially decaying trend of the ROS as fuel moisture increases (live fuels) and the nearly linearly decaying trend of the ROS as relative humidity increases (dead fuels). Moreover, the model reproduces the increasing ROS when the surface coverage or fuel load increases, showing a threshold for no-fire conditions.
@@
A limitation of this model is that the effect of radiation is only considered locally. The model uses the Rosseland approximation to model the local radiative effects as a diffusion mechanism. Therefore, heat is transferred between adjacent areas in a manner similar to conduction and the long-distance effect of radiation is not represented, which results in short pre-heating times. Consequently, the influence of topography must be approximated as an advective phenomenon. The present approach offers a good balance between accuracy and complexity in the model; nevertheless, it can be enhanced by means of additional physics.
@@
So far, the optical path length $\delta$ was treated as a constant. A more detailed modeling approach may include a connection between the optical path length and the vegetation structure, for example, via the surface coverage $SC$, the fuel fraction $Y$ or the radiation parameter $k$. Something else to explore in the future is the dependence of the wind correction coefficient, $\beta$, on the moisture content, which may improve the model’s predictions in presence of strong wind conditions.
@@
Accounting for fuel moisture effects on wildfire dynamics opens up possibilities to consider hydrological and plant physiological controls on the wildfire propagation. This is already done in the context of wildfire risk, see, for example, [47, 48], but is not common for wildfire propagation. Another interesting topic that could be explored with this model is the linkage between vegetation pattern dynamics in water-limited systems and wildfire propagation in these systems. A spatially aggregated modeling approach has been applied in [18] to study wildfire regime–vegetation feedback. In this context, a spatially distributed modeling of wildfire propagation could further provide insights into wildfire–ecosystem structure relations.

### v2

@@ 4.5 Two-dimensional domains
Now the spatial domain is extended to two spatial dimensions (2D) and different scenarios with spatial heterogeneity are considered. Below, results are presented for four different cases and the observations made above for one spatial dimension are now extended to more realistic fire propagation patterns in two-dimensional landscapes.
@@ 4.5.1 Propagation in a homogeneous environment: assessing the effect of fuel moisture
The aim of this case is to assess the effect of fuel moisture in 2D fire propagation across homogeneous land, under a low wind velocity $\mathbf{w}=(0.25,0.25)^{T}$ m/s. The computational domain is $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
with $(x_{1},y_{1})=(100,100)$ m. Under these wind conditions, the predominant propagation direction points toward the top-right corner of the domain.
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 1400 seconds using $\mathsf{CFL}=0.4$. The complete model with radiation and constant reaction rate is used, assuming the simplification in Eq. (9). The S2M model is used to represent the effect of moisture. The model parameters are taken from Tab. 24, unless otherwise specified.
Four different simulations are compared considering different moisture levels, $M=0.0$, $M=0.2$, $M=0.4$, $M=0.6$. Fig. 14 shows the predicted temperature and biomass distribution at the final time. Besides, the fire isochrones are depicted in red every 200 seconds. These results evidence that the fuel moisture level has a strong impact on the shape of the burned area, as well as on the rate of spread. When considering zero moisture, the propagation takes place in all directions, being slightly faster in the direction of wind. Contrarily, when increasing the moisture level, propagation mainly takes place in the direction of wind. Besides, the width of the fire front and the rate of spread reduces as the moisture level increases. From a geometric perspective, the angle of the fire front decreases with increasing moisture level.
@@
Figure 14: Fire spread for different moisture content levels with a wind velocity $\mathbf{w}=(0.25,0.25)^{T}$ m/s, starting at $(x,y)=(100,100)$. Temperature and remaining fuel mass fraction are shown at the final time. The red isochrones provide information on the temporal evolution of the fire. Contour levels are depicted from 300 K to 1400 K.
The reduction of the rate of spread in direction of the wind is in accordance with Fig. 10 showing the decrease in spreading rate with increasing moisture level. More detailed, the spread orthogonal to the wind direction is reduced the most. This is in line with the zero wind case of Fig. 10 where the rate of spread drops to zero for moisture levels above 80%. In contrast, for non-zero wind speed, the decrease in the rate of spread is less significant and reaches a non-zero threshold. The two-dimensional simulation combines these two results when regarding either the direction of the wind or the orthogonal direction. As the moisture levels regarded are all below the critical threshold of $M=0.8$, there is still a fire spread orthogonal to the wind velocity, but the rate of spread in this direction is small.
Tab. 7 shows the total burned area at the final time of the simulation. The burned area is calculated considering the points with a biomass fraction below the threshold $Y=0.99$. The results show a strong reduction of burned area for higher moisture levels, e.g. a reduction of 80% for a moisture level $M=0.2$ compared to the no moisture case.
@@ 4.5.2 Propagation in heterogeneous environments: assessing the effect of fuel moisture
The aim of this case is to understand the model behavior when changing the fuel moisture in a heterogeneous landscape composed of four quadrants with different fuel moisture content. The computational domain is $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 1400 seconds using $\mathsf{CFL}=0.4$. The complete model with radiation and constant reaction rate is used, assuming the simplification in Eq. (9). The S2M model is used to represent the effect of moisture. The model parameters are taken from Tabs. 24, unless otherwise specified, with zero wind conditions.
@@
Figure 15: Temperature (left) and biomass fraction (right) for fire propagation in a case with a different moisture content in each quadrant. The isochrones show continuous connections of the solution at the intersections of different moisture content regions. Contour levels are depicted from 300 K to 1400 K.
@@ 4.5.3 Propagation in heterogeneous environments: assessing the effect of surface coverage
The aim of this case is to understand the model behavior when changing the fuel load by modifying the surface coverage, $SC$, assuming that the fuel load is a limiting factor for the parameters chosen here –i.e., corresponding to the region of Fig 12 where the ROS increases with fuel load–. Those areas with a lower value of $SC$ are expected to act as a firebreak. The computational domain is $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 2200 seconds using $\mathsf{CFL}=0.4$. The complete model with radiation and constant reaction rate is used, assuming the simplification in Eq. (9). The S2M model is used to represent the effect of moisture. The model parameters are taken from Tabs. 24, unless otherwise specified above.
Four different simulations consider different fuel surface coverages, $SC_{r}$, namely $SC_{r}=0.2$, $SC_{r}=0.4$, $SC_{r}=0.6$, $SC_{r}=0.8$, setting a wind velocity $\mathbf{w}=(1,1)^{T}$ m/s. Fig. 16 shows the predicted temperature and biomass distribution at the final time. Besides, the fire isochrones are depicted in red every 200 seconds. These results confirm that higher fuel loads, regarded as higher $SC$, enhance fire propagation. Besides, there is a threshold at which the fire extinguishes and propagation stops, acting as a firebreak. In this example, the threshold appears for a value of $SC$ between 0.2 and 0.3. Fig. 16 shows a strong dependency on stopping the fire on the surface coverage but, in oppose to the influence of the moisture content, the rate of spread is less strongly affected by the reduced surface coverage. This tendency is supported by the one-dimensional simulation in Fig. 13, where the drop from a higher rate of spread for high surface coverage values to zero for small surface coverage values is very abrupt.
@@
Figure 16: Fire spread at the interface of a changing surface coverage ($SC$), starting from $(50,50)$ m with a wind velocity $\mathbf{w}=(0.25,0.25)^{T}$ m/s. The temperature and remaining fuel mass fraction are shown at final time, the isochrones give information on the temporal evolution. The area with $SC=0.2$ acts as a fire break, areas with low $SC$ slow down the rate of spread. Contour levels are depicted from 300 K to 1400 K.
@@ 4.5.4 Propagation in heterogeneous environments with topography
This case examines fire propagation across uneven vegetation, both with and without the influence of topography. The computational domain is $[0,200]\times[0,200]$ m2. The initial condition for the temperature is given by
@@
The computational mesh has 400 cells in each Cartesian direction. The simulation runs for 2600 seconds using $\mathsf{CFL}=0.4$. The complete model with radiation and constant reaction rate is used, assuming the simplification in Eq. (9). The bulk density and heat capacity are computed using the more realistic mixture approximation in Eq. (8). The moisture is zero and the wind is given by $\mathbf{w}=(0.0,0.5)^{T}$ m/s. The model parameters are taken from Tables 24, unless otherwise specified above.
Fig. 17 shows the evolution of the fire (isocrhones) and the final state for the temperature and fuel load. A tendency to spread to areas with higher fuel load is observed. Additionally, there are fire breaks for small fuel loads, which is comparable to the situation in Fig. 16(a). The wind direction is parallel to the $y$-axis, but there is a strong propagation as well in the areas with high fuel load orthogonal to the wind direction.
Fig. 17 shows the spread of fire in the same biomass configuration but now with additional uneven topography. The topography acts as an additional advective influence, compare Sec. 2.3. The topography in Eq. (87) has a gradient pointing towards $(x,y)=(200,200)$ m, therefore there is an additional advective influence towards the upper right-hand corner. Fig. 17 shows this influence by having a stronger tendency to spread towards this corner.
@@
Figure 17: Temperature (left) and fuel load (right) at the final time for the simulation of fire with (b) and without topography (a), considering spatially heterogeneous fuel with wind $\mathbf{w}=(0.0,0.5)^{T}$ m/s, $M=0.0$, and the mixture approximation in Eq. (8). (a) Flat terrain is considered: The fire spreads faster in areas with high fuel load, influenced by the wind in $y-$direction. (b) Gaussian topography is considered: The slope towards $(x,y)=(200,200)$ enhances the fire spread in this direction.
4.5.5 Validation using an experimental field study: case F19
The experimental case F19 [35] belongs to a set of experiments conducted in the Northern Territory of Australia, during July and August 1986 [17, 18]. For all the experiments, fuels were open grasslands and grasses were continuous and normally fully cured; besides, the topography was flat. For case F19, the fuel was Themeda grass. The grass-land plot was 200 × 200 m and the ignition line fire was 175 m long. This line fire was created with drip torches carried by two field workers walking for 56 s in opposite directions from the center point to the ends of the line fire. The average wind speed was 4.8 m/s. The resulting average rate of spread is around 1.45 m/s (87 m/min).
The experiment F19 has been extensively used in the literature for model validation [35] and is herein reproduced using the proposed model for the same purpose. A spatial domain $[0,250]\times[0,250]$ m2 is considered, setting the remaining biomass fraction at zero along a 25 m wide strip along the edges of the domain. The resolution of the mesh is 1 m, which leads to an average time step of 0.05 s. The initial condition is set as a constant temperature –ambient temperature– in the whole domain. To reproduce the ignition process, a moving heat source term is added on the left side of the domain, mimicking the operation carried out by the field workers.
The parameter values for this case are shown in Table 8. It must be noted that there are different choices for the parameters reproducing the observed ROS, with differences in the width of the fire front, maximal temperature, and other variables. Furthermore, note that the ROS in this case is unusually high, which amplifies the difference between the parameters in Table 8 and the default parameters. A domain height of $l=1$ m is considered, but other values will be explored in the future to obtain more suitable parameter values, e.g. for the optical path length.
Figure 18 shows the fire front at $t=56$ s, $t=86$ s and $t=138$ s reproduced by the model and it is compared with the experimental measurements from [18, 17] and the numerical results provided by the WFDS model from [35]. The proposed model reproduces the reported ROS and the shape of the fire front, which evolves from a triangular shape to a more rounded shape. The computed results closely match those of the WFDS model [35].
Table 8: Parameter values for the case F19, adapted from [35].
Parameter Symbol Value Units
Ground elevation $Z$ 0 m
Wind velocity $\mathbf{w}$ 4.8 m·s-1
Initial fuel volume ratio $R_{f,0}$ 0.00061 -
Fuel density $\bar{\rho}_{f}$ 512 kg·m-3
Initial fuel load $W_{f,0}$ 0.31 kg·m-2
Specific heat of dry fuel $c_{p,f0}$ 1.5 kJ·kg-1·K-1
Fuel pyrolysis temperature $T_{pc}$ 400 K
Fuel combustion heat $\mathcal{H}$ 15000 kJ·kg-1
Reaction rate $A_{c}$ 0.1 s-1
Fuel moisture content $M$ 0.06 -
Thermal conductivity $k_{c}$ 0.0001 kW·m-1· K-1
Optical path length $\delta$ 10.0 m
Ambient cooling coefficient $\hat{\alpha}$ 0.2 kW·m-2·K-1
Wind correction coefficient $\beta$ 0.15 -
Figure 18: Fire fronts at $t=56$ s, $t=86$ s and $t=138$ s reproduced by the model (shading and contour lines), compared with the experimental measurements from [17, 18] (black markers) and the numerical results provided by the WFDS model from [35] (green dashed line). Contour lines are shown for the temperature every 100 K, from 900 K to 1400 K. Temperatures below 900 K are represented in light yellow (background color) to better visualize the fire front.
@@ 5 Conclusions
We presented an ADR wildfire propagation model that accounts for the effect of fuel moisture, and also considers wind, local radiation, natural convection, and topography. Special attention has been given to the derivation of the governing equations from the fundamental conservation laws within an idealized vegetation stratum which is modelled as a two-phase porous flow. The proposed model includes different mechanisms for live and dead fuel by using the apparent calorific capacity method, which models the effect of thermal phase changes through an apparent heat capacity. Three different moisture models–the S2M, C2M, and S3M–have been presented to represent the evaporation process and the constituents of fuel with different levels of detail, being the S2M the preferred option when total fuel moisture data are available and the C2M model used when willing to distinguish between live and dead fuels. The use of the C2M model will benefit wildfire propagation predictions, because dead fuel moisture is often considered to be the main control on fire characteristics and fire impact on fuel and vegetation, see, e.g., [44, 63]. A limitation of this model is that the relative humidity does not vary over time. A more complete approach, see, e.g., [41], could be easily adapted. Also note that the utilization of dead fuel moisture data in the model, if available, is straightforward by substituting this approach –i.e., $M_{d}$ would be considered as an input parameter–.
Our work provides new insights on ADR-based wildfire propagation modeling that complement previous studies, e.g., [7, 28, 30, 37, 40, 45, 68]. Key modeling decisions have been carefully analyzed and justified through both theoretical analysis and simulation results. One of the fundamental findings of this analysis is related to the degree of simplification that can be applied to the terms in the advection-diffusion-reaction equation. Whereas the combustion heat and atmospheric cooling source terms can be significantly simplified without excessively compromising the physical realism of the solution, the diffusion term requires, at least, a non-linear model for the diffusion parameter to obtain physically consistent results. This is addressed through Rosseland’s approximation of local radiation. With regard to the combustion heat source term, it is shown that the commonly used Arrhenius combustion function can be approximated by a constant as long as the switching function for activating and deactivating the combustion process is still included. This approximation is in line with the macroscopic scale of the wildfire model: The Arrhenius law is based on chemical processes on a much smaller length scale, and the resulting non-linear combustion term contrasts the model accuracy on the macroscopic scale. For the combustion model, the key parameters are the combustion heat of the fuel, $\mathcal{H}$, and the kinetic constant, $A_{c}$, the latter depending on the physio-chemical structure of the fuel. With respect to radiation, the fundamental aspect when considering local radiation through Rosseland’s approximation lies in the strong non-linearity of the diffusion term. The dependence of the diffusion mechanism on $T^{3}$ links the ROS to the energy of the medium, which is proportional to its temperature, ensuring physically consistent results and reproducing the observed trends. The optical path length, $\delta$, is the key parameter for the Rosseland approximation and can be retained as a calibration parameter.
@@
The parametric analyses presented herein provide evidence of the model’s ability to reproduce the expected fire propagation behavior and the dependence of the ROS on the main drivers of fire spread. The model reproduces both the exponentially decaying trend of the ROS as fuel moisture increases (live fuels) and the nearly linearly decaying trend of the ROS as relative humidity increases (dead fuels). Moreover, the model reproduces an increase in ROS with the fuel load when the latter is a limiting factor, showing a threshold for no-fire conditions. When the fuel load is not a limiting factor, a Ricker-like function behavior is reproduced, with an exponentially decaying ROS for large fuel load values.
@@
A limitation of this model is that the effect of radiation is only considered locally. The model uses the Rosseland approximation to model the local radiative effects as a diffusion mechanism. Therefore, heat is transferred between adjacent areas in a manner similar to conduction and the long-distance effect of radiation is not represented, which results in short pre-heating times. Consequently, the influence of topography must be approximated as an advective phenomenon. The use of non-local radiation models would avoid this unphysical representation of topography and would improve the effect of wind, which should not be only represented through the advective term. In any case, the present approach based on local radiation offers a good balance between accuracy and complexity in the model.
@@
So far, the optical path length $\delta$ was treated as a constant. A more detailed modeling approach may include a connection between the optical path length and the vegetation structure, for example, through the packing ratio $R_{f}$, the fuel fraction $Y$, or the radiation parameter $k$. Something else to explore in the future is the dependence of the wind correction coefficient, $\beta$, on the moisture content, which may improve the model’s predictions in presence of strong wind conditions. Additionally, a more complete definition of the chemical reaction processes through the consideration of oxygen, water vapor, and combustion products would also improve the physical realism of the model.
@@
Accounting for fuel moisture effects on wildfire dynamics opens possibilities to consider hydrological and plant physiological controls on the wildfire propagation. This is already done in the context of wildfire risk, see, for example, [50, 51], but is not common for wildfire propagation. Another interesting topic that could be explored with this model is the link between vegetation pattern dynamics in water-limited systems and wildfire propagation in these systems. A spatially aggregated modeling approach has been applied in [20] to study wildfire regime–vegetation feedback. In this context, spatially distributed modeling of wildfire propagation could further provide insights into wildfire–ecosystem structure relations.